OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. Slope 3/4 through (-8,4) The equation of the line is y=

OpenStudy (anonymous):

@SolomonZelman Another? :)

OpenStudy (solomonzelman):

y - y1 = m (x - x1) plug in your numbers, and I;ll correct you if anything... okay?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

How do I plug the numbers in? @SolomonZelman

OpenStudy (anonymous):

Straight-line equations, or "linear" equations, graph as straight lines, and have simple variable expressions with no exponents on them. If you see an equation with only x and y — as opposed to, say x2 or sqrt(y) — then you're dealing with a straight-line equation. There are different types of "standard" formats for straight lines; the particular "standard" format your book refers to may differ from that used in some other books. (There is, ironically, no standard definition of "standard form".) The various "standard" forms are often holdovers from a few centuries ago, when mathematicians couldn't handle very complicated equations, so they tended to obsess about the simple cases. Nowadays, you likely needn't worry too much about the "standard" forms; this lesson will only cover the more-helpful forms. thier is some help

OpenStudy (solomonzelman):

y - y1 = m (x - x1) (-8,4) y - 4 = 3/4 (x - (-8) )

OpenStudy (solomonzelman):

Can you finish the y - 4 = 3/4 (x - (-8) ) ?

OpenStudy (anonymous):

Then solve @SolomonZelman ?

OpenStudy (solomonzelman):

yes. I would rather use the word "simplify" ... expand the parenthesis on the right side, and then isolate the y by adding 4 to both sides.

OpenStudy (anonymous):

y=3/4 x +4 @SolomonZelman ?

OpenStudy (solomonzelman):

close.

OpenStudy (solomonzelman):

y - 4 = 3/4 (x - (-8) ) y - 4 = 3/4 (x + 8) y - 4 = 3/4x + (3/4)*8 showing how to multiply 3/4 times 8.... y - 4 = 3/4x + (3*8 /4) y - 4 = 3/4x + (24/4) y - 4 = 3/4x + 6 y - 4 + 4 = 3/4x + 6 + 4 y = 3/4x + 10